# Hypothesis-Driven Theory-of-Mind Reasoning for Large Language Models

### v1

@@ 3.2 Comparison with Reasoning Models
Thought-tracing applied models show better performance than reasoning models.
Except for o1 on MMToM-QA, models with thought-tracing significantly outperform o1, o1-mini, o3-mini, DeepSeek R1, and QwQ 32B preview on all benchmarks. [^1] Reasoning models were provided with extra formatting prompts to align with the evaluation setup of these benchmarks. Without those formatting prompts, the scores are significantly lower. Moreover, even LLMs with CoT performs better than reasoning models, such as o3-mini, DeepSeek R1 and QwQ 32B. The o1 model also underperforms than CoT on FANToM. This suggest that reinforcing mathematical or programming reasoning does not generalize to social reasoning, necessitating separate training scheme in this domain.
@@
A performance trade-off can be observed between false belief scenarios and true belief scenarios in Paraphrased ToMi. While o1 and o3-mini achieve near-perfect scores (i.e., 100) on the false belief scenarios, their performance on the easier true belief scenarios drops sharply to below 50 and even below 30 in some cases. A similar pattern is observed with DeepSeek R1, whose scores on the easier true belief scenarios are lower than those on the false belief scenarios. In contrast, the QwQ 32B preview follows the pattern seen in other vanilla LLMs, displaying higher scores for true belief scenarios but lower scores for false belief scenarios. Although these reasoning models do not release their training data, an interesting future direction would be to investigate how reinforcement learning influences this trade-off. On the other hand, thought-tracing applied models show a more balanced performance across false belief and true belief scenarios, with improvements observed in both cases, except for GPT-4o.
Figure 2: The average output token counts of reasoning models and thought-tracing for Paraphrased ToMi and MMToM-QA.
@@
Figure 2 summarizes the average output token counts for o3-mini, o1, DeepSeek R1, and GPT-4o with thought-tracing on Paraphrased ToMi.
@@
(3) Token usage for incorrect and correct responses to ToM questions differs across datasets. Although a recent analysis reported incorrect responses from DeepSeek R1 and QwQ-32B preview have significantly higher token usage (Wang et al., 2025), this does not hold for ToM benchmarks. In Paraphrased ToMi, all reasoning models show significantly higher token usage for incorrect ToM responses than correct ones. Moreover, the pattern is different for MMToM-QA, which the token usage is similar between correct and incorrect responses or is higher for correct ones. On the other hand, thought-tracing exhibits balanced token usage regardless of whether the ToM responses are correct or incorrect. We note that the relative difficulty of the questions in ToMi and MMToM-QA is consistent compared to benchmarks in other domains, and all models achieve near-perfect scores on ToMi’s factual questions.
(4) Models with thought-tracing produce significantly shorter reasoning trajectories while achieving better performance. As shown in Table 2, GPT-4o+TT and GPT-4o+TT+CoT deliver the best and second-best performance among these models on ToMi, with scores of 76 and 74.5, respectively. Despite the better performance, these models generate significantly fewer output tokens than o3-mini, o1 (with medium and high reasoning effort), and DeepSeek R1. Furthermore, compared to o1-low and o3-mini-low, GPT-4o+TT and GPT-4o+TT+CoT score almost 10 points higher. Similarly in MMToM-QA, GPT-4o+TT+CoT outperforms DeepSeek R1 and shows performance comparable with o3-mini-medium while using less than 50% of the tokens.
@@
Additionally, we analyze the thought switching patterns for DeepSeek R1, following (Wang et al., 2025), on Paraphrased ToMi and MMToM-QA. The authors initially reported that frequent thought switching leads to underthinking and ultimately lower performance. For Paraphrased ToMi, incorrect responses included an average of 9.69 thoughts, while correct ones included an average of 6.21 thoughts, which is consistent with the reported pattern. However, for MMToM-QA, responses for incorrect responses included an average of 6.83 thoughts similar to correct responses, which contained an average of 6.45 thoughts.
3.3 Analysis on Thought-tracing
@@
We manually evaluated thought-tracing’s errors on MMToM-QA and find an inherent bias across models. In many cases, when an agent searches for something and repeatedly encounters an object—despite not interacting with it—the models generate hypotheses suggesting that the object is what the agent is looking for. However, according to both social common-sense and explicit models of ToM (Baker et al., 2017; Ying et al., 2024), if an agent repeatedly sees an object but continues searching elsewhere, it is likely that they are actually searching for something else. Since this error may be a side effect of the perception inference in thought-tracing, we conduct ablation experiments.
@@
Table 3: Score difference comparing variants of thought-tracing to the original thought-tracing on Paraphrased ToMi, FANToM, and MMToM-QA.
@@
We also tested another variant of thought-tracing that self-corrects its previous hypothesis and predict a new hypothesis during the propagation step (§2.2). Table 3 shows the results. Although self-correction has shown performance increase in some domains (Madaan et al., 2023; Shinn et al., 2023), it does not show better performance for hypotheses propagation on the three ToM benchmarks. In most cases, it leads to worse performance, indicating self-correction capability is not effective in our iterative updates, potentially due to the accumulation of correction errors. We suggest future works to investigate the effects of LLM’s self-correcting behavior in the social domain.
@@ 4 Related Work
Cognitive science research has shown that humans reason about other agents’ minds by assuming their actions are rational and goal-directed (Dennett, 1981; Csibra et al., 1999; Baillargeon et al., 2016). These processes have been formalized as Bayesian Theory of Mind (Baker et al., 2017; Jara-Ettinger et al., 2019), which models agents as rational actors that plan and act upon their beliefs and desires to achieve their goals. This probabilistic model can then be inverted via inverse planning, inferring agents’ mental states from their behavior (Baker et al., 2009; Ying et al., 2023; 2024), including via SMC methods (Zhi-Xuan et al., 2020; 2024a). While thought-tracing is patterned on this model-based approach, we do not explicitly model rational agents, but instead use LLMs as implicit models of agents’ behavior, and as hypothesis generators for agents’ thoughts, enabling our method to be applied to open-ended inputs.
@@
has emerged as a pivotal area of research, particularly in mathematics, coding, and puzzle solving. Early work, such as Chain-of-Thought (Wei et al., 2022), demonstrated the benefits of generating intermediate reasoning steps, and was later augmented by approaches capable of exploring multiple reasoning paths (Yao et al., 2023; Besta et al., 2024), aggregation across reasoning chains (Wang et al., 2023), and problem decomposition (Zhou et al., 2023). Recently, reasoning models trained via reinforcement learning—o1 (OpenAI et al., 2024b), R1 (DeepSeek-AI et al., 2025a), and QwQ (Qwen-Team, 2024)—have shown remarkable performance. Most of these approaches leverage objectively verifiable answers to enable reinforcement and accurate selection among multiple reasoning paths. However, such verification is challenging in social reasoning due to subjectivity and uncertainty. To handle uncertainty, recent methods perform probabilistic inference in LLMs via sequential Monte Carlo (Lew et al., 2023b; Zhao et al., 2024; Loula et al., 2025), but have not been used to infer and track mental states. In this work, we introduce a general SMC-like algorithm capable of operating effectively in the social domain while bypassing the need for objective verification.
@@ 5 Conclusion
In this paper, we introduced thought-tracing, a new inference-time algorithm that performs approximate inference over agents’ mental states. Our approach draws inspiration from the sequential Monte Carlo (SMC) method and follows the conceptual structure of Bayesian Theory of Mind (BToM; Baker et al., 2017). Thought-tracing leverages an SMC-like structure by using LLMs as proposals for natural language hypotheses, which are then weighted with actions from the given text—all without requiring ground-truth answers to questions in datasets. Our evaluation on four theory-of-mind benchmarks demonstrates its effectiveness, significantly improving upon baseline LLMs and outperforming recent reasoning models such as o3-mini and R1. Additionally, our findings reveal that reasoning models show different behaviors in the social domain compared to their remarkable performance in areas like math and coding. Nonetheless, there remains room for improvement due to errors in areas such as how hypotheses about agents’ goals are generated (Section 3.3). This suggests that more work needs to be done to achieve the reliability that explicit BToM models demonstrate in closed-world settings (Zhi-Xuan et al., 2022; 2024b; Ying et al., 2024). We hope these insights pave the way for future research on inference-time reasoning in social settings, where AI agents will play an increasingly important role.
@@ Appendix A BigToM Re-annotation
Backward Belief (FB) Backward Belief (TB) Forward Action (FB) Forward Action (TB) Forward Belief (FB) Forward Belief (TB)
@@
Appendix B Thought-tracing Setup
@@
We use greedy decoding through out our thought-tracing algorithm. We set the number of hypotheses to 4 across all experiments. The threshold for the effective sample size is set to 2. We measure the text similarity across hypotheses using the Jaccard similarity, and the threshold for rejuvenation is set to 0.25. For action labeling, we use GPT-4o-2024-08-06 for ParaphrasedToMi and BigToM, and a rule-based algorithm for FANToM and MMToM-QA.

### v2

@@ 3.2 Comparison with Reasoning Models
ThoughtTracing applied models show better performance than reasoning models.
Except for o1 on MMToM-QA, models with ThoughtTracing significantly outperform o1, o1-mini, o3-mini, DeepSeek R1, and QwQ 32B preview on all benchmarks. [^2] Reasoning models were provided with extra formatting prompts to align with the evaluation setup of these benchmarks. Without those formatting prompts, the scores are significantly lower. Moreover, even LLMs with CoT perform better than reasoning models, such as o3-mini, DeepSeek R1 and QwQ 32B. The o1 model also underperforms than CoT on FANToM. This suggests that reinforcing mathematical or programming reasoning does not generalize to social reasoning, necessitating separate training scheme in this domain.
@@
A performance trade-off can be observed between false belief scenarios and true belief scenarios in ParaphrasedToMi. While o1 and o3-mini achieve near-perfect scores (i.e., 100) on the false belief scenarios, their performance on the easier true belief scenarios drops sharply to below 50 and even below 30 in some cases. A similar pattern is observed with DeepSeek R1, whose scores on the easier true belief scenarios are lower than those on the false belief scenarios. In contrast, the QwQ 32B preview follows the pattern seen in other vanilla LLMs, displaying higher scores for true belief scenarios but lower scores for false belief scenarios. Notably, QwQ significantly underperforms on FANToM, primarily due to excessive false positives. For instance, it frequently predicts that characters are aware of certain information when they are, in fact, unaware. This issue arises consistently in both list-type and binary (yes/no) questions.
Although these models do not release their training data, an interesting future direction would be to investigate how reinforcement learning influences this trade-off. On the other hand, ThoughtTracing applied models show more balanced performance across false belief and true belief scenarios, with improvements observed in both cases, except for GPT-4o.
@@
Figure 2 summarizes the average output token counts for o3-mini, o1, DeepSeek R1, and GPT-4o with ThoughtTracing on ParaphrasedToMi.
@@
(3) Token usage for incorrect and correct responses to ToM questions differs across datasets. Although a recent analysis reported incorrect responses from DeepSeek R1 and QwQ-32B preview have significantly higher token usage (Wang et al., 2025), this does not hold for ToM benchmarks. In ParaphrasedToMi, all reasoning models show significantly higher token usage for incorrect responses to ToM questions than correct ones. Moreover, the pattern is different for MMToM-QA, which the token usage is similar between correct and incorrect responses or is higher for correct ones. On the other hand, ThoughtTracing exhibits balanced token usage regardless of whether the responses to the ToM questions are correct or incorrect. We note that the relative difficulty of the questions in ToMi and MMToM-QA is consistent compared to benchmarks in other domains, and all models achieve near-perfect scores on ToMi’s factual questions.
(4) Models with ThoughtTracing produce significantly shorter reasoning traces while achieving better performance. As shown in Table 2, GPT-4o+TT and GPT-4o+TT+CoT deliver the best and second-best performance among these models on ToMi, with scores of 76 and 74.5, respectively. Despite better performance, these models generate significantly fewer output tokens than o3-mini, o1 (with medium and high reasoning effort), and R1. Furthermore, compared to o1-low and o3-mini-low, GPT-4o+TT and GPT-4o+TT+CoT score almost 10 points higher. Similarly in MMToM-QA, GPT-4o+TT+CoT outperforms R1 and shows performance comparable with o3-mini-medium while using less than 50% of the tokens.
Figure 2: The average number of tokens in the reasoning trace from the reasoning models and ThoughtTracing for ParaphrasedToMi and MMToM-QA.
@@
We also analyze the thought switching patterns for DeepSeek R1, following (Wang et al., 2025), on ParaphrasedToMi and MMToM-QA. The authors initially reported that frequent thought switching leads to underthinking and ultimately lower performance. For ParaphrasedToMi, incorrect responses included an average of 9.69 thoughts, while correct ones included an average of 6.21—consistent with the reported pattern. However, for MMToM-QA, responses for incorrect responses included an average of 6.83 thoughts similar to correct responses, which contained an average of 6.45.
3.3 Analysis on ThoughtTracing
@@
We manually evaluated ThoughtTracing’s errors on ParaphrasedToMi, FANToM, and MMToM-QA. Due to the model’s near-perfect performance on BigToM, we did not conduct error analysis, as the number of errors was too limited to draw meaningful conclusions.
(1) ParaphrasedToMi: We observed a recurring pattern of incorrect perception predictions. In true belief scenarios, the model sometimes incorrectly infers that the agent did not witness another agent moving an object—often due to a conservative estimation of perception (e.g., assuming the agent was not paying attention).
(2) FANToM: The model tends to overestimate the target agent’s prior knowledge, occasionally assuming familiarity among characters. However, FANToM is designed such that characters are meeting for the first time, making these inferences incorrect.
These cases from ParaphrasedToMi and FANToM reflect our intentional design choice to avoid benchmark-specific assumptions—such as presuming agents always observe everything or have predefined relationships. We believe this is a principled trade-off for generalizability. Incorporating benchmark-specific cues or more explicit assumptions in the prompts will reduce these errors.
(3) MMToM-QA: We find an inherent bias across models. In many cases, when an agent searches for something and repeatedly encounters an object—despite not interacting with it—the models generate hypotheses suggesting that the object is what the agent is looking for. However, according to both social common sense and explicit models of ToM (Baker et al., 2017; Ying et al., 2025), if an agent repeatedly sees an object but continues searching elsewhere, it is likely that they are actually searching for something else. Since this error may be a side effect of perception inference in ThoughtTracing, we conduct ablation experiments.
@@
Table 3: Score difference comparing variants of ThoughtTracing to the original algorithm on ParaphrasedToMi, FANToM, and MMToM-QA.
@@
We also tested another variant of ThoughtTracing that self-corrects its previous hypothesis and predict a new hypothesis during the propagation step (§2.2). Table 3 shows the results. Although self-correction has shown a performance increase in some domains (Madaan et al., 2023; Shinn et al., 2023), it does not show better performance for hypotheses propagation on the three ToM benchmarks. In most cases, it leads to worse performance, indicating self-correction capability is not effective in our iterative updates, potentially due to the accumulation of correction errors. We suggest future works to investigate the effects of LLM’s self-correcting behavior in the social domain.
@@ 4 Related Work
Cognitive science research has shown that humans reason about other agents’ minds by assuming their actions are rational and goal-directed (Dennett, 1981; Csibra et al., 1999; Baillargeon et al., 2016). These processes have been formalized as Bayesian Theory of Mind (Baker et al., 2017; Jara-Ettinger et al., 2019), which models agents as rational actors that plan and act upon their beliefs and desires to achieve their goals. This probabilistic model can then be inverted via inverse planning, inferring agents’ mental states from their behavior (Baker et al., 2009; Ying et al., 2023; 2025), including via SMC methods (Zhi-Xuan et al., 2020; 2024a). While ThoughtTracing is patterned on this model-based approach, we do not explicitly model rational agents, but instead use LLMs as implicit models of agents’ behavior, and as hypothesis generators for agents’ thoughts, enabling our method to be applied to open-ended inputs.
@@
Inference-time reasoning has emerged as a pivotal area of research for LLMs, particularly in mathematics, coding, and puzzle solving. Early work, such as Chain-of-Thought (Wei et al., 2022), demonstrated the benefits of generating intermediate reasoning steps, and was later augmented by approaches capable of exploring multiple reasoning paths (Yao et al., 2023; Besta et al., 2024), aggregation across reasoning chains (Wang et al., 2023), and problem decomposition (Zhou et al., 2023). Recently, reasoning models trained via reinforcement learning—o1 (OpenAI, 2024b), R1 (DeepSeek-AI, 2025a), and QwQ (Qwen-Team, 2024)—have shown remarkable performance. Most of these approaches leverage objectively verifiable answers to enable reinforcement and accurate selection among multiple reasoning paths. However, such verification is challenging in social reasoning due to subjectivity and uncertainty. To handle uncertainty, recent methods perform probabilistic inference in LLMs via sequential Monte Carlo (Lew et al., 2023b; Zhao et al., 2024; Loula et al., 2025), but have not been used to infer and track mental states. In this work, we introduce a general SMC-like algorithm capable of operating effectively in the social domain while bypassing the need for objective verification.
@@ 5 Conclusion
In this paper, we introduced ThoughtTracing, a new inference-time algorithm that performs approximate inference over agents’ mental states. Our approach draws inspiration from the sequential Monte Carlo (SMC) method and follows the conceptual structure of Bayesian Theory of Mind (BToM; Baker et al., 2017). ThoughtTracing leverages an SMC-like structure by using LLMs as proposals for natural language hypotheses, which are then weighted with actions from the given text—all without requiring ground-truth answers to questions in datasets. Our evaluation on four theory-of-mind benchmarks demonstrates its effectiveness, significantly improving upon baseline LLMs and outperforming recent reasoning models such as o3-mini and R1. Additionally, our findings reveal that reasoning models show different behaviors in the social domain compared to their remarkable performance in areas like math and coding. Nonetheless, there remains room for improvement due to errors in areas such as how hypotheses about agents’ goals are generated (Section 3.3). This suggests that more work needs to be done to achieve the reliability that explicit BToM models demonstrate in closed-world settings (Zhi-Xuan et al., 2022; 2024b; Ying et al., 2025). We hope these insights pave the way for future research on inference-time reasoning in social settings, where AI agents will play an increasingly important role.
Acknowledgments
This work was supported in part by the Amazon Research Gift Grant. Tan Zhi-Xuan is funded by the Open Philanthropy AI Fellowship.
@@ Appendix A BigToM Re-annotation
Backward
Belief (FB)
Backward
Belief (TB)
Forward
Action (FB)
Forward
Action (TB)
Forward
Belief (FB)
Forward
Belief (TB)
@@
Appendix B ThoughtTracing Setup
@@
We use greedy decoding throughout our ThoughtTracing algorithm. We set the number of hypotheses to 4 across all experiments. The threshold for the effective sample size is set to 2. We measure the text similarity across hypotheses using the Jaccard similarity, and the threshold for rejuvenation is set to 0.25. For action labeling, we use GPT-4o-2024-08-06 for ParaphrasedToMi and BigToM, and a rule-based algorithm for FANToM and MMToM-QA. For FANToM, we cache the traced thoughts of each character in the conversation for efficiency.
Appendix C Input Prompts used in ThoughtTracing
For generalization across different models and benchmarks, we use simple prompts for ThoughtTracing, leaving more sophisticated prompts for future work. Note, the additional tokens introduced by ThoughtTracing are bounded by the length of the input text.
Perception Inference You are an expert perception tracker tasked with determining whether {target agent} perceived the target context. Briefly describe what {target agent} saw or why {target agent} could not see the target context. Make your response concise. {context history} <target context>{current context}</target context>
Initialization {context} Generate a numbered list of {n} hypotheses what were {target agent}’s thoughts (e.g., beliefs, intent) that led to the action above. Do not add any additional comments.
Propagation You are an expert assistant trying to predict {target agent}’s thoughts. <previous context> {context} </previous context> <previous prediction regarding {target agent}’s thoughts> {hypothesis} </previous prediction regarding {target agent}’s thoughts> <current context> {new context} </current context> What did {target agent} believe?
Weight Update Your job is to evaluate the probability of actions/utterance under a given fact. Use common sense: for instance, if someone is searching for an item, they are likely to take it once they find it rather than merely observing it. If they don’t take it and just sees it, it indicates a lack of interest and that was not what they were looking for. Briefly explain the probability of the action/utterance under the given fact first and then give the answer option with prefix ’Answer:’ {context} <{target agent}’s thoughts> {hypothesis} </{target agent}’s thoughts> <next action>{action}</next action> <note>{perception}</note> Question: Based on the context and {target agent}’s thoughts provided, would {target agent} do the next actions or say the next utterances described above? Let’s think step by step and give the final answer. (a) Very Likely (Around 90%) (b) Likely (Around 70%) (c) Somewhat Likely (Around 60%) (d) Somewhat Unlikely (Around 25%) (e) Unlikely (Around 20%) (f) Very Unlikely (Below 10%)
Rejuvenation Your task is to paraphrase the following text. Make sure to keep the meaning of the text intact while rephrasing them. Do not add any additional comments. {hypothesis}
Hypotheses Summarization {context} <{target agent}’s thoughts> Prediction 1 (Weight: {weight 1}) {hypothesis 1} Prediction 2 (Weight: {weight 2}) {hypothesis 2} Prediction N (Weight: {weight N}) {hypothesis N} </{target agent}’s thoughts> Question: What did {target agent} believe?
Appendix D Example Thought Tracing Process
Given the length of the interleaved hypotheses, we provide a summarized example of how the hypotheses evolve over time and how the weights are updated during the tracing process in Table 5.
State 1: …In the Room B, there is a sofa, a cabinet, a desk, and a coffee table. …The Room D is furnished with a sofa, a coffee table, and a desk. …The Room C is equipped with eight cabinets, a microwave, a fridge, a dishwasher, a Room C table, and a stove. …Jessica is situated in Room B.
Action 1: She proceeds towards the Room A, approaches the Room A cabinet, opens it, and subsequently closes it.
Hypotheses: (Weight: 0.33) Jessica was curious about the contents of the cabinet in Room A, driven by a general exploratory behavior. (Weight: 0.33) Jessica had a specific task or goal that required her to verify the contents of the cabinet in Room A, leading her to open and then close it after finding it empty. (Weight: 0.20) Jessica believed that Room A might contain items of interest or value, motivating her to investigate the cabinet there. (Weight: 0.12) Jessica intended to check if the cabinet in Room A was similar to the one in Room B, possibly looking for specific objects.
State 2: None
Action 2: She then navigates her way to the Room C, heads towards the fourth Room C cabinet, opens it, and then shuts it.
Hypotheses: (Weight: 0.25) Jessica was curious and exploratory; she noted the variability in cabinet contents (empty in Room A, Object K in Room C), leading her to believe contents differ across rooms. (Weight: 0.25) Jessica had a specific goal, believed cabinets might hold a desired item, and methodically searched until she found Object K in Room C, reinforcing her belief in a purposeful search. (Weight: 0.25) Jessica believed cabinets might contain valuable items, remained curious despite an empty cabinet in Room A, and was validated by finding Object K in Room C. (Weight: 0.25) Jessica expected cabinets in other rooms to contain objects like those in Room B; after an empty cabinet in Room A, she kept searching and discovered Object K in Room C, altering her expectations.
Table 5: A summarized example of the ThoughtTracing process on MMToM-QA.
The first set of hypotheses is initialized from State 1 and Action 1. We update their weights using Action 1. Since the diversity of the hypotheses is above the threshold, rejuvenation is not performed. Additionally, the effective sample size is above the threshold, so resampling is not carried out.
Next, we derive the second set of hypotheses by propagating from the first set. Consequently, the second set inherits the contents of its parent hypotheses. The likelihood of the second set is uniform, as all four previous hypotheses are approximately equally effective in predicting that Jessica is heading to Room C (i.e., Action 2).
Appendix E Experiment Results
Table 6 shows the full evaluation results on ParaphrasedToMi, FANToM, and MMToM-QA.
Model ParaphrasedToMi FANToM MMToM-QA First-order False Belief Second-order False Belief First-order True Belief Second-order True Belief Belief Multi-choice Info Accessibility List type Info Accessibility Yes/No type Answerability List type Answerability Yes/No type Type 1.1 Type 1.2 Type 1.3 Type 2.1 Type 2.2 Type 2.3 Type 2.4 GPT-4o 15.0 62.0 98.0 63.0 64.5 38.5 91.8 63.0 79.7 94.0 40.7 67.9 17.7 50.0 1.6 56.7 GPT-4o + CoT 69.0 83.0 79.0 37.0 83.9 78.8 93.7 74.2 83.4 100.0 41.2 81.2 24.1 64.0 4.2 55.0 GPT-4o + TT 89.0 80.0 85.0 50.0 76.1 82.4 96.2 83.0 92.4 100.0 91.2 56.2 10.3 76.0 8.3 55.0 GPT-4o + TT + CoT 94.0 72.0 79.0 52.0 89.1 80.4 97.3 77.4 92.9 100.0 94.1 84.4 41.4 68.0 0.0 65.0 o1-preview 100.0 100.0 57.0 15.0 83.9 69.2 91.9 81.5 91.4 94 98.8 90.1 35.5 95.0 3.3 52.2 o1-low-effort 98.0 99.0 58.0 16.0 67.4 52.9 95.1 66.0 83.0 94.4 100.0 96.9 82.8 72.0 0.0 80.0 o1-medium-effort 98.0 94.0 64.0 27.0 67.4 60.8 95.1 60.4 87.0 88.9 100.0 96.9 79.3 100.0 4.2 75.0 o1-high-effort 99.0 95.0 64.0 34.0 70.6 63.0 92.4 62.1 84.9 88.9 100.0 96.9 86.2 100.0 0.0 75.0 o1-mini 46.0 74.0 79.0 40.0 65.2 39.2 84.8 67.9 82.9 100.0 85.3 87.5 0.0 68.0 0.0 50.0 o3-mini-low-effort 55.0 69.0 78.0 53.0 34.8 19.6 80.3 35.8 71.3 100.0 44.1 68.8 20.7 80.0 0.0 55.0 o3-mini-medium-effort 87.0 94.0 61.0 16.0 34.8 23.5 84.6 43.4 76.7 100.0 88.2 93.8 24.1 92.0 0.0 60.0 o3-mini-high-effort 99.0 100.0 52.0 5.0 30.4 41.2 87.3 54.7 78.1 100.0 100.0 90.6 31.0 88.0 0.0 65.0 DeepSeek R1 62.0 100.0 86.0 35.0 86.3 66.7 89.5 75.9 88.2 97.2 44.1 78.1 6.9 56.0 0.0 40.0 Llama 3.3 70B 17.0 72.0 94.0 46.0 69.9 0.0 83.5 18.5 76.1 100.0 20.6 25.0 55.2 56.0 0.0 60.0 Llama 3.3 70B + CoT 53.0 80.0 79.0 46.0 73.9 27.5 90.2 52.8 84.6 100.0 17.6 25.0 58.6 60.0 0.0 60.0 Llama 3.3 70B + TT 68.0 56.0 95.0 66.0 63.0 27.5 90.0 52.8 87.3 100.0 38.2 37.5 3.4 52.0 0.0 65.0 Llama 3.3 70B + TT + CoT 88.0 87.0 83.0 50.0 71.7 35.3 96.2 56.6 91.9 100.0 94.1 40.6 20.7 60.0 0.0 70.0 Gemini 1.5 Pro 26.0 61.0 83.0 49.0 71.7 2.0 81.3 18.9 79.9 100.0 47.1 62.5 0.0 72.0 0.0 50.0 Gemini 1.5 Pro + CoT 61.0 80.0 71.0 21.0 80.4 29.4 82.7 52.8 80.8 100.0 52.9 62.5 0.0 72.0 0.0 50.0 Gemini 1.5 Pro + TT 80.0 86.0 85.0 46.0 80.4 80.4 86.2 86.8 89.0 100.0 82.4 71.9 0.0 64.0 0.0 25.0 Gemini 1.5 Pro + TT + CoT 87.0 60.0 79.0 20.0 82.6 76.5 89.3 79.2 91.3 94.4 88.2 56.2 10.3 84.0 29.2 75.0 Qwen 2.5 72B 22.0 56.0 85.0 66.0 34.8 5.9 87.3 34.0 49.6 97.2 14.7 37.5 10.3 60.0 12.5 55.0 Qwen 2.5 72B + CoT 50.0 76.0 84.0 61.0 76.1 37.3 90.8 49.1 43.6 97.2 14.7 37.5 10.3 60.0 12.5 50.0 Qwen 2.5 72B + TT 65.0 61.0 91.0 80.0 48.9 21.6 95.2 52.8 84.6 100.0 47.1 31.2 6.9 64.0 8.3 45.0 Qwen 2.5 72B + TT + CoT 90.0 70.0 86.0 58.0 70.7 68.6 96.8 71.7 90.5 94.4 61.8 31.2 41.4 56.0 4.2 65.0 QwQ 32B Preview 28.0 27.0 89.0 91.0 67.4 0.0 59.5 18.9 55.9 97.2 61.8 34.4 3.4 52.0 16.7 55.0
Table 6: Full evaluation results across three theory-of-mind benchmarks. The notation ‘+$T$’ denotes that our ThoughtTracing method has been applied, while ‘+CoT’ indicates that chain-of-thought has been used. We exclude BigToM because most of the models achieved a near-perfect score on our re-annotated subset.